### v1

@@ Abstract
I Introduction
@@
We attempt a training of an RL agent in a custom Gymnasium environment. The goal is to develop a strategy to autonomously lock a Fabry-Perot (FP) optical cavity operating in a non-linear regime using a deep reinforcement learning (RL) agent, for future applications in GW detectors [5]. The work began with the design and validation of a highly accurate time-domain simulator that models cavity dynamics under varying mirror velocities, including cavity ring-down effects at high finesse. This simulator is embedded into a custom Gymnasium environment, enabling the RL agent to interact with realistic cavity conditions.
@@
II Improvements to the actor-critic architecture
@@
Further, we discussed a possible evolution of our system toward meta reinforcement learning (Meta-RL) [13]. This approach would shift the DRL algorithm design problem to another level. Instead of manually tuning the algorithm, we would incorporate the tuning process into training in order to achieve better generalization, adaptability and efficiency across a wide spectrum of tasks.
Summarizing possible strategies, we think that currently Meta-RL training is computationally too heavy for us and that the risk of producing algorithms unstable in implementation is too high. We therefore concluded that our current system will benefit from a temporal unit ( recurrent neural network (RNN) or gated recurrent unit (GRU)). Adding such a layer helps extract information from order-sensitive sequential data. We will focus on this development first.
III Low-latency execution

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Phys. Proc.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
We attempt a training of an RL agent in a custom Gymnasium [5] environment. The goal is to develop a strategy to autonomously lock a Fabry-Perot (FP) optical cavity operating in a non-linear regime using a deep reinforcement learning (RL) agent, for future applications in GW detectors [6]. The work began with the design and validation of a highly accurate time-domain simulator that models cavity dynamics under varying mirror velocities, including cavity ring-down effects at high finesse. This simulator is embedded into a custom Gymnasium environment, enabling the RL agent to interact with realistic cavity conditions.
@@
2 Current status
Trained agent is able to lock a cavity at the first resonance. See an example of a locking run in Fig. 1. The sequence starts at a random cavity length out of resonance. At the beginning, the agent changes the cavity length with the maximal allowed speed and when it finds the resonance, it modifies mirror movements accordingly to its policy. We would like to stress that the environment used for this training and demonstration does not simulate the finite force exerted by real actuators; thus, it does not simulate the inertia of the suspended mirror.
In such an environment, the agent training lasts from two to three minutes. At the first glance, it looks quick; however, we noticed that the training time does not change significantly when executed on two distinctly different hardware platforms. On both platforms, the environment calculates with floating point precisions. The first platform is an M1 chip (8 core CPU, 8 core GPU, RAM ) reaching realistically [8] while the second is a CUDA accelerated NVIDIA RTX A2000 GPU ( VRAM) reaching almost (considering 80% of nominal computing power). Clearly, the problem lays in the environment implementation, which does not benefit from parallel execution. Such slow training does not allow efficient domain randomization and further obstructs meta-learning training. We are going to address these issues in the next software release, together with optimization of the most time-consuming procedures using Numba [9].
3 Improvements to the actor-critic architecture
@@
Further, we discussed a possible evolution of our system toward meta reinforcement learning (Meta-RL) [16]. This approach would shift the DRL algorithm design problem to another level. Instead of manually tuning the algorithm, we would incorporate the tuning process into training in order to achieve better generalization, adaptability and efficiency across a wide spectrum of tasks. At the moment, we do not have many diverse tasks to justify a continuous optimization of the policy leveraging meta-learning. However, modifying the environment in the future in order to support parallelization will enable the exploitation of the more robust generalization approach offered by meta-learning, compared to domain randomization. Summarizing possible strategies, we think that currently Meta-RL training is computationally too heavy for us and that the risk of producing algorithms unstable in implementation is too high. We therefore concluded that our current system will benefit from a temporal unit ( recurrent neural network (RNN) or gated recurrent unit (GRU)). Adding such a layer helps extract information from order-sensitive sequential data. We will focus on this development first.
4 Low-latency execution